OpenStudy (anonymous):

i need help celeste wants to have her hair cut and perm and also go to lunch. she knows she will need $68. the perm costs twice as much as her hair cut and she needs $5 for lunch. how much does the perm costs

OpenStudy (anonymous):

Let x be the cost of hair cut so now the cost of Perm will be twice of x which is 2x and cost of lunch is $5 and total cost for all the three is $68. U can just now put them together and u get the answer

OpenStudy (anonymous):

x + 2x + 5 = 68 3x = 63 x = 63/3 = $ 21 Cost of Hair cut is 21$ and cost of perm is 2x which is 2(21) = $42

OpenStudy (anonymous):

thanks